Three charges +4q,Q,+q are placed in a straight line with Q in the midway betweenthe other two charges.If net force on +q is zero,then Q must be

Grade:12

2 Answers

Arun
4 years ago

two force are acting on q ,one due to 4q and second due to Q

 

F1 = 1/4∏ε (4q)q/l2 

F2 = 1/4∏ε (Q)q/(l/2)2 

      = 1/4∏ε (4Q)q/l2 

 F1 +F2 =0

1/4∏ε (4q)q/l2   +  1/4∏ε (4Q)q/l2  =0

 

 Q=-q

Khimraj
4 years ago

F1 = 1/4∏ε (4q)q/l2 

F2 = 1/4∏ε (Q)q/(l/2)2 

      = 1/4∏ε (4Q)q/l2 

 F1 +F2 =0

1/4∏ε (4q)q/l2   +  1/4∏ε (4Q)q/l2  =0

 

 Q=-q